12. a) Create an example of a quadratic
equation that can be factored and
solved with integer solutions.
b) Create an example of a quadratic
equation that can be factored and
solved with non-integer solutions.

Answer by lynnlo(4176)

(1)x^2-x=6=====================problem
x^2-x-6=0====move terms to one side
(x-3)(x+2)=0======factor
(x-3)=0 OR (x+2)=0=====solve
x=3 OR X=-2=========SOLUTION
(2)√2x-3+1=4==========problem
√2x=6
x=3√2
x=18=========solve by factoring